# Democratizing Agentic AI with Fast Test-Time Scaling on the Edge

### v1

@@ 6.2. End-to-End Performance Improvement
Completion Latency. Beyond improving goodput, FlashTTS also delivers substantial reductions in end-to-end completion latency. Fig. 13 shows that FlashTTS achieves an average latency reduction of 38% to 68% across all configurations and n values compared to the vLLM baseline.
The latency breakdown within FlashTTS reveals the distinct performance characteristics of each model configuration. In the 7B+1.5B configuration, generator latency (unfilled portion) is the dominant cost. Conversely, in the 1.5B+7B configuration, the larger 7B verifier model contributes a substantial portion of the total latency, becoming nearly on par with the generator as n increases.
FlashTTS effectively reduces both the generation and verification components of latency. On average, it reduces verifier latency by 75% to 85% and reduces generation latency by 36% to 66% across all n values. The dramatic reduction in verifier latency is primarily driven by our LookAhead Verification technique, which enhances computational locality by pre-verifying tokens. The substantial decrease in generator latency is achieved through the combined effects of our other optimizations: Asymmetric Multi-Model Memory Allocation and Dynamic Prefix-Aware Scheduling enhance KV cache efficiency, while Speculative Beam Extension hides straggler latency by utilizing idle GPU cycles. We note one exception where verifier latency slightly increases (at n=512 for the 1.5B+7B model on AMC), a direct trade-off from our memory allocator prioritizing the heavily-loaded generator by reducing the verifier’s KV cache capacity.
@@ 6.3. Algorithm Performance
While FlashTTS is designed to guarantee algorithmic equivalence with the baseline, minor variations in output can occur since our scheduling optimizations may alter the sampling order. We now analyze these effects.
Top-1 Accuracy. As shown in Fig. 14(a), the Top-1 accuracy of FlashTTS is highly competitive with the baseline. On the more challenging AIME dataset, FlashTTS consistently matches or slightly improves accuracy, likely because its speculative execution focuses computation on the more promising reasoning paths. In general, both perform comparably, confirming the algorithm equivalence.
@@ 6.4.2. In-depth Study of Speculative Beam Extension
As shown in Fig. 16 (left), Speculative Beam Extension improves hardware utilization. While the baseline vLLM implementation suffers from progressively decaying GPU compute utilization as faster reasoning paths in a batch finish early, FlashTTS maintains a higher and more consistent utilization by speculatively generating tokens for completed beams. The overall latency is also reduced, as the speculative tokens generated in one iteration can be used as a head start for the next, shortening their required generation time. The performance of Speculative Beam Extension is also affected by its truncation ratio, R. As shown in Fig. 16 (right), a higher ratio (R=0.85), which aggressively retains speculative work, yields more goodput improvement.
@@ 7.1. Reasoning Systems and Speculative Execution
Edge and Memory-Constrained Serving: The recent development of edge LLM serving systems focuses mainly on optimizing the deployment of non-reasoning workloads (Zheng et al., 2025; Song et al., 2024; Yu et al., 2024; Wei et al., 2025). Although Certaindex (Fu et al., 2024b) covers LLM reasoning serving, it focuses solely on CoT reasoning with query-level scheduling and early termination. It does not handle the irregular computation patterns within TTS reasoning trajectories, nor does it optimize scheduling between the generator and verifier.
Speculative Execution: The philosophy of speculative generation for LLMs was initially explored at the algorithmic level, primarily through speculative and parallel decoding techniques (Hu et al., 2025a; Li et al., 2024, 2025; Qin et al., 2025; Cheng, [n. d.]; Fu et al., 2024a; Cai et al., 2024; Chen et al., 2024). However, these methods serve as algorithmic enhancements aimed at accelerating LLM decoding by enabling multi-token generation. Speculative generation has also been applied in retrieval-augmented generation (RAG) systems (Jin et al., 2024; Zhang et al., 2024; Hu et al., 2025b), where it is used to prefetch or cache retrieved documents to reduce latency.
@@ 7.2. Memory Management and Prefix Sharing
Beyond the various memory and scheduling optimizations (Kwon et al., 2023; Zhong et al., 2024; Wu et al., 2024a; Sun et al., 2024; Lee et al., 2024; Fu et al., 2024c; Lin et al., 2024; Kim et al., 2025a), two other methods have emerged for improving memory efficiency in LLM serving: offloading and prefix sharing.
Offloading for LLMs: Offloading is a primary strategy for alleviating LLM memory pressure, typically via computation offloading and data offloading. In computation offloading, several approaches, such as FlexGen (Sheng et al., 2023), FastDecode (He and Zhai, 2024), PowerInfer (Song et al., 2024), and NEO (Jiang et al., 2024), distribute LLM pipelines across CPU and GPU to reduce GPU load. LIA (Kim et al., 2025b) explores offloading to Intel AMX, while (Yuan et al., 2025) supports edge–cloud partitioning for inference. In data offloading, DeepSpeed-Inference (Aminabadi et al., 2022) utilizes CPU host memory for activation offloading, while (Alizadeh et al., 2024) explores flash-based offloading strategies. AIF (Lee et al., 2025) pushes this further by supporting in-flash processing to reduce data movement overhead.
Prefix Caching and Sharing: Most prior work on prefix caching and sharing focuses on query-level optimization using tree-structured management (Kwon et al., 2023; Zheng et al., 2024b). BatchLLM (Zheng et al., 2024a) explores the global prefix reuse through ahead-of-time prefix identification. FastTree (Pan et al., 2025a) improves tree-structured inference via context-query grouping to enhance cache locality. RAGCache (Jin et al., 2024) investigates prefix sharing with dynamic overlapping between the retrieval and inference steps for retrieval-augmented generation. KVFlow (Pan et al., 2025b) further advances prefix scheduling in multi-agent systems by introducing workflow-aware eviction policies and overlapped KV prefetching. However, previous work on prefix optimization has primarily focused on coarse-grained, query-level sharing in non-reasoning LLM serving scenarios. In contrast, LLM reasoning workloads introduce new opportunities for fine-grained, prefix-aware sharing during decoding.
@@ 8. Conclusion
This paper presents FlashTTS, a plug-and-play third-party serving system that makes Test-Time Scaling (TTS) both practical and efficient on memory-constrained edge devices. By analyzing the common execution pattern of mainstream TTS methods, we identify several key system challenges: i) hardware underutilization of irregular reasoning search paths, ii) suboptimal cache reuse, and iii) memory pressure from multi-model execution. FlashTTS addresses these challenges through three novel synergistic techniques: Speculative Beam Extension, Dynamic Prefix-Aware Scheduling, and Asymmetric Multi-Model Memory Allocation. Our evaluation shows that FlashTTS enables low-latency, high-quality reasoning using edge LLMs for memory-constrained devices; it narrows their performance gap with cloud-scale models, and advances the vision of democratized agentic AI.

### v2

@@ 6.2. End-to-End Performance Improvement
Completion Latency. Beyond improving goodput, FastTTS also delivers substantial reductions in end-to-end completion latency. Fig. 13 shows that FastTTS achieves an average latency reduction of 38% to 68% across all configurations and n values compared to the vLLM baseline.
The latency breakdown within FastTTS reveals the distinct performance characteristics of each model configuration. In the 7B+1.5B configuration, generator latency (unfilled portion) is the dominant cost. Conversely, in the 1.5B+7B configuration, the larger 7B verifier model contributes a substantial portion of the total latency, becoming nearly on par with the generator as n increases.
FastTTS effectively reduces both the generation and verification components of latency. On average, it reduces verifier latency by 75% to 85% and reduces generation latency by 36% to 66% across all n values. The dramatic reduction in verifier latency is primarily driven by our LookAhead Verification technique, which enhances computational locality by pre-verifying tokens. The substantial decrease in generator latency is achieved through the combined effects of our other optimizations: Asymmetric Multi-Model Memory Allocation and Dynamic Prefix-Aware Scheduling enhance KV cache efficiency, while Speculative Beam Extension hides straggler latency by utilizing idle GPU cycles. We note one exception where verifier latency slightly increases (at n=512 for the 1.5B+7B model on AMC), a direct trade-off from our memory allocator prioritizing the heavily-loaded generator by reducing the verifier’s KV cache capacity.
@@ 6.3. Algorithm Performance
While FastTTS is designed to guarantee algorithmic equivalence with the baseline, minor variations in output can occur since our scheduling optimizations may alter the sampling order. We now analyze these effects.
Top-1 Accuracy. As shown in Fig. 14(a), the Top-1 accuracy of FastTTS is highly competitive with the baseline. On the more challenging AIME dataset, FastTTS consistently matches or slightly improves accuracy, likely because its speculative execution focuses computation on the more promising reasoning paths. In general, both perform comparably, confirming the algorithm equivalence.
@@
Figure 15. Goodput improvements on constrained hardware and coding tasks. Experiments are conducted on NVIDIA RTX 3070 Ti (8GB) and RTX 4070 Ti (12GB) GPUs using the AIME dataset, and on the HumanEval code generation benchmark using an RTX 4090. Offloading is used on the RTX 3070 Ti; as a result, lower absolute goodput is observed.
6.4. Generality on Hardware and Benchmarks
To demonstrate the generality of FastTTS across more resource-constrained GPUs and diverse tasks, we further extend our evaluation to additional devices and code generation workloads.
[…]
@@ 6.5.2. In-depth Study of Speculative Beam Extension
As shown in Fig. 17 (left), Speculative Beam Extension improves hardware utilization. While the baseline vLLM implementation suffers from progressively decaying GPU compute utilization as faster reasoning paths in a batch finish early, FastTTS maintains a higher and more consistent utilization by speculatively generating tokens for completed beams. The overall latency is also reduced, as the speculative tokens generated in one iteration can be used as a head start for the next, shortening their required generation time. The performance of Speculative Beam Extension is also affected by its truncation ratio, R. As shown in Fig. 17 (right), a higher ratio (R=0.85), which aggressively retains speculative work, yields more goodput improvement.
@@ 7.1. Reasoning Systems and Speculative Execution
Edge and Memory-Constrained Serving: The recent development of edge LLM serving systems focuses mainly on optimizing the deployment of non-reasoning workloads (Zheng et al., 2025; Song et al., 2024; Yu et al., 2024; Wei et al., 2025).
Efficient Reasoning: Although Certaindex (Fu et al., 2024c) covers LLM reasoning serving, it focuses solely on CoT reasoning with query-level scheduling and early termination. It does not handle the irregular computation patterns within TTS reasoning trajectories, nor does it optimize scheduling between the generator and verifier. Beyond system-level serving optimizations, recent algorithmic advances such as speculative reasoning (Pan et al., 2025a; Fu et al., 2025; Liao et al., 2025) aim to accelerate inference by leveraging an efficient draft model. However, these approaches modify the output distribution and therefore lack algorithmic equivalence. We emphasize that such algorithmic techniques are orthogonal to our system-level optimizations and can be seamlessly integrated into our framework to achieve further speedups.
Speculative Execution: The philosophy of speculative generation for LLMs was initially explored at the algorithmic level, primarily through speculative and parallel decoding techniques (Hu et al., 2025a; Li et al., 2024, 2025; Qin et al., 2025; Cheng, Ph.D. Thesis; Fu et al., 2024b; Cai et al., 2024; Chen et al., 2024). However, these methods serve as algorithmic enhancements aimed at accelerating LLM decoding by enabling multi-token generation. Speculative generation has also been applied in retrieval-augmented generation (RAG) systems (Jin et al., 2024; Zhang et al., 2024; Hu et al., 2025b), where it is used to prefetch or cache retrieved documents to reduce latency. In contrast, our Speculative Beam Extension targets the unique system-level bottleneck of parallel reasoning: hardware underutilization caused by irregular step lengths. Instead of verifying draft tokens, we utilize idle compute slots from completed paths to speculatively extend future reasoning steps, thereby mitigating straggler effects and maintaining high GPU occupancy without altering the algorithm.
@@ 7.2. Memory Management and Prefix Sharing
Beyond the various memory and scheduling optimizations (Kwon et al., 2023; Zhong et al., 2024; Wu et al., 2024a; Sun et al., 2024; Lee et al., 2024; Fu et al., 2024a; Lin et al., 2024; Kim et al., 2025b), two other methods have emerged for improving memory efficiency in LLM serving: offloading and prefix sharing.
Offloading for LLMs: Offloading is a primary strategy for alleviating LLM memory pressure, typically via computation offloading and data offloading. In computation offloading, several approaches, such as FlexGen (Sheng et al., 2023), FastDecode (He and Zhai, 2024), PowerInfer (Song et al., 2024), and NEO (Jiang et al., 2024), distribute LLM pipelines across CPU and GPU to reduce GPU load. LIA (Kim et al., 2025a) explores offloading to Intel AMX, while (Yuan et al., 2025) supports edge–cloud partitioning for inference. In data offloading, DeepSpeed-Inference (Aminabadi et al., 2022) utilizes CPU host memory for activation offloading, while (Alizadeh et al., 2024) explores flash-based offloading strategies. AIF (Lee et al., 2025) pushes this further by supporting in-flash processing to reduce data movement overhead.
Prefix Caching and Sharing: Most prior work on prefix caching and sharing focuses on query-level optimization using tree-structured management (Kwon et al., 2023; Zheng et al., 2024a). BatchLLM (Zheng et al., 2024b) explores the global prefix reuse through ahead-of-time prefix identification. FastTree (Pan et al., 2025b) improves tree-structured inference via context-query grouping to enhance cache locality. RAGCache (Jin et al., 2024) investigates prefix sharing with dynamic overlapping between the retrieval and inference steps for retrieval-augmented generation. KVFlow (Pan et al., 2025c) further advances prefix scheduling in multi-agent systems by introducing workflow-aware eviction policies and overlapped KV prefetching. However, previous work on prefix optimization has primarily focused on coarse-grained, query-level sharing in non-reasoning LLM serving scenarios. In contrast, LLM reasoning workloads introduce new opportunities for fine-grained, prefix-aware sharing during decoding.
@@ 8. Conclusion
This paper presents FastTTS, a plug-and-play third-party serving system that makes Test-Time Scaling both practical and efficient on memory-constrained edge devices. By analyzing the common execution pattern of mainstream TTS methods, we identify several key system challenges: i) hardware underutilization of irregular reasoning search paths, ii) suboptimal cache reuse, and iii) memory pressure from multi-model execution. FastTTS addresses these challenges through three novel synergistic techniques: Speculative Beam Extension, Dynamic Prefix-Aware Scheduling, and Asymmetric Multi-Model Memory Allocation. Our evaluation shows that FastTTS enables low-latency, high-quality reasoning using edge LLMs for memory-constrained devices. It narrows their performance gap with cloud-scale models, and advances the vision of democratized agentic AI.
9. Acknowledgement
The support of UKRI and EPSRC (grant numbers UKRI256, EP/V028251/1, EP/N031768/1, EP/S030069/1, and EP/X036006/1), KIAT, AMD and Intel is gratefully acknowledged.
@@ A.2. Proof of Local Optimality
Appendix B Artifact Appendix
B.1. Abstract
FastTTS is a framework for fast test-time scaling on edge devices, leveraging speculative beam extension, prefix caching, and adaptive memory allocation. The artifact includes the complete source code, benchmark scripts, and configuration files to reproduce the goodput, latency, and accuracy results presented in the paper. It supports evaluating AIME 2024 and AMC 2023 datasets using automated workflows.
B.2. Artifact check-list (meta-information)
  • Algorithm: FastTTS
  • Program: Python 3.11
  • Model: Qwen2.5-Math-1.5B-Instruct, Qwen2.5-Math-7B-Instruct, Math-Shepherd-Mistral-7B-PRM, Skywork-o1-Open-PRM-Qwen-2.5-1.5B
  • Data set: AIME 2024, AMC 2023 (via HuggingFace)
  • Run-time environment: Linux, Conda, Python 3.11, CUDA 12.9, PyTorch 2.7.0
  • Hardware: NVIDIA GeForce RTX 4090 GPU (24 GB VRAM)
  • Execution: Automated via run_all_experiments.py
  • Metrics: Goodput, Latency, Accuracy
  • Output: JSONL logs, PDF plots
  • Experiments: Verifier-guided Beam Search across datasets to reproduce goodput, latency, accuracy results.
  • How much disk space required (approximately)?: $\sim$200GB (including model weights and environment).
  • How much time is needed to prepare workflow (approximately)?: $\sim$15 minutes
  • How much time is needed to complete experiments (approximately)?: $\sim$60 GPU hours (depending on GPU)
  • Publicly available?: No
  • Data licenses (if publicly available)?: MIT/Apache 2.0 (Dataset dependent)
  • Workflow automation framework used?: Custom Python scripts
B.3. Description
B.3.1. How to access
The artifact is available here.
B.3.2. Hardware dependencies
The experiments require a Linux machine with at least one NVIDIA GeForce RTX 4090 GPU (24 GB VRAM).
B.3.3. Software dependencies
Described in the code.
B.3.4. Data sets
The artifact uses the following datasets, automatically downloaded via the Hugging Face Datasets library:
  • AIME 2024
  • AMC 2023
B.3.5. Models
The following models are automatically downloaded from Hugging Face Hub:
  • Generators:
    • Qwen/Qwen2.5-Math-1.5B-Instruct
    • Qwen/Qwen2.5-Math-7B-Instruct
  • Verifiers:
    • peiyi9979/math-shepherd-mistral-7b-prm
    • Skywork/Skywork-o1-Open-PRM-Qwen-2.5-1.5B
B.4. Installation
For details, please read the README.md file in the repository.
  1. (1)
    Download the repository.
  2. (2)
    Create and activate the Conda environment:
    conda env create -f environment.yml
    conda activate FastTTS
  3. (3)
    Install the package in editable mode:
    pip install -e .
  4. (4)
    Install the modified Skywork inference module:
    cd modified-skywork-o1-prm-inference
    pip install -e .
    cd ..
B.5. Experiment workflow
The main results can be reproduced using:
run_all_experiments.py
[…]
python run_all_experiments.py --exp
This command runs the benchmark configurations defined in benchmarks/configs/.
To generate the plots from the collected results:
python run_all_experiments.py --plot
To run both experiments and plotting in sequence:
python run_all_experiments.py --exp --plot
B.6. Evaluation and expected results
Upon successful execution, the script generates the following PDF figures in benchmarks/benchmark_results/figs/:
  • main_results_combined.pdf: Compares Goodput across different methods and model sizes. It corresponds to Figure 12.
  • latency_combined.pdf: Shows the latency breakdown. It corresponds to Figure 13.
  • acc.pdf: Displays the accuracy metrics for the evaluated benchmarks. It corresponds to Figure 14(a).
The results should match the trends reported in Figure 12, Figure 13, and Figure 14a of the paper, demonstrating the efficiency improvements.
B.7. Experiment customization
Experiments can be customized by modifying the YAML configuration files located in benchmarks/configs/. Individual benchmarks can be run via benchmarks/run_benchmarks.py with a specific config file.
B.8. Notes
Ensure that the conda environment is activated before running any scripts. The initial run may take longer due to model downloads.